OpenStudy (anonymous):

Anyone here good at f(x) to know if they are odd or even..... f(x)=5x^6-7/x^3

OpenStudy (anonymous):

So, something is even if f(x) = f(-x). Therefore, to test if the function is even: \[f(x) = 5x^6 - 7/x^3 \neq 5x^6 + 7/x^3 = 5(-x)^6 - 7/(-x)^3 = f(-x)\] Since f(x) does not equal f(-x), the function is not even. If something is odd, then f(x) = -1 * f(x) Therefore, to test if the function is odd: \[f(x) = 5x^6 - 7/x^3 \neq -5x^6 - 7/x^3 = -5(-x)^6 + 7/(-x)^3 = -1 * f(-x)\] Since f(x) does not equal -1 * f(-x) the function is not odd.

OpenStudy (anonymous):

wow, the tutorial is wanting me to substitute (x) with (-x) and work it out...